OpenStudy (anonymous):

Analyze the graph of y= x^3 -3x +2 by finding the required information.

OpenStudy (cwrw238):

first find y' can you do that?

OpenStudy (anonymous):

Yes, I already found them all, but some of them got wrong so I am correcting, I just don't which ones specifically I got wrong. Can you help with that? my y- intercepts I got (0,2)

OpenStudy (cwrw238):

yes thats correct for y intecept y = 2 when x = 0

OpenStudy (cwrw238):

y' = 3x^2 - 3 3x^2 - 3 = 0 at turning points on the graph 3x^2 = 3 x^2 = 1 and x = +1 , -1 so you have truning points at x = 1 and x = -1

OpenStudy (cwrw238):

do you know how to find the nature of these turning points?

OpenStudy (anonymous):

Ok, that one I got wrong, I had 0,1 . Then I have no vertical or horizontal asymptotes. For increasing intervals I have ( -1, 0) , decreasing intervals ( 0,3) I think I got that one wrong. Then relative maxima (0,2), and relative minima (1,0) ?

OpenStudy (cwrw238):

do they want you to determine the max and min by finding increasing / decreasing intervals? y

OpenStudy (anonymous):

Well, I first found the relatives max and min, then the increasing and decreasing intervals, followed by the inflection points and finally the convape up and down intervals... however, I think I got the increasing and decreasing int. completely wrong?

OpenStudy (cwrw238):

darn it i just lost my post (how i dont know)

OpenStudy (cwrw238):

relative maximums is at (-1,4) and relative minimum is at (1,0)

OpenStudy (cwrw238):

ignore my 3rd last post - I messed up!! I've tried to delete it but the system wont let me.

OpenStudy (anonymous):

Yeah, hate it when that happens :) Oh, ok, so the max I got wrong. Uh, my concave up intervals are ( - infinity, - 1/2) , ( 1/2, +infinity) ; cocave down ( -1/2, 1/2) . My inflection points are 6,0

OpenStudy (anonymous):

ok

OpenStudy (cwrw238):

geez i must admit i've forgotten how to do those concave up and down turning points

OpenStudy (cwrw238):

have you found the zeroes of the function or dont they want those?

OpenStudy (cwrw238):

as there is a minimum at the point (1,0) that is also a zero of the function

OpenStudy (anonymous):

No, I haven't found the zeros

OpenStudy (cwrw238):

x^3 - 3x + 2 = 0 there must bea xero < -1 where there's a maximum try -2: (-2)^3 - 3(-2) + 2 = 0 ok theres a zero at (-2,0) and and at (1,0) as I explained above

OpenStudy (cwrw238):

the graph will lokk a bit like